# Easy-access online social media metrics can effectively identify misinformation sharing users

### v1

@@ Discussion
In this study, we aimed to provide a low-barrier approach to distinguishing between individuals who are more and less likely to share misinformation on social media. Drawing upon existing literature and leveraging easily accessible data from X, we tested four social network metrics as potential predictors of user factuality. Our analysis focused on follower count, average daily tweet count, number of users followed, and account age. The results offer several insights into the effectiveness of these metrics as indicators of misinformation sharing tendencies.
Our findings indicate that when considered individually, each of the four metrics—follower count, average daily tweet count, number of users followed, and account age—has a significant relationship with user factuality scores. However, when all metrics were analysed together using multinomial regression, the distinctive power of follower count and followed account count diminished for the separation of low and high factuality, while average daily tweet count and number of users followed remained robust predictors of low and high factuality scores.
@@
Finally, the multinomial regression analysis revealed certain significant interaction effects in the case of follower and followed account count. Among those who follow a low number of accounts, less popular users are more likely to be highly factual, while among those who follow many accounts, popular users are more likely to be highly factual. Similarly, among those with a low number of average daily tweets, less popular users are more likely to be highly factual, while among those who tweet a lot, more popular users are more likely to be highly factual. These cross-effects highlight the context-dependent nature of the relationship between these social network metrics and factuality.
@@
These findings have several important implications. For social media platforms, the results suggest that basic metrics like tweet count and account age could be used to identify users more likely to share misinformation, allowing for targeted interventions such as accuracy prompts or digital literacy campaigns. Policymakers could advocate for the incorporation of these metrics into misinformation mitigation strategies, ensuring a low-cost, scalable approach. Future research should continue to explore additional accessible data points and refine existing metrics to enhance the accuracy of identifying misinformation-prone users.
In conclusion, this study demonstrates that basic, easily accessible social network metrics, particularly tweet count and account age, can effectively differentiate between low and high factuality users. These studied metrics provide a practical and efficient method for identifying users susceptible to sharing misinformation, offering a foundation for cost-effective interventions to curb the spread of misinformation on social media platforms.
@@ Methods
Our study builds upon prior studies by Rathje et al. [18] and (Galeazzi et al. 2023), and to carry out the analysis we rely on the X (Twitter at the time) dataset collected by these authors. Rathje et al. [18] gathered 463 survey responses through the online recruitment platform Prolific. Participants were asked to respond to survey questions and requested to provide their X handles. Building on this, Galeazzi et al. (2023) obtained the usernames of all those Prolific participants who provided their X handles, then collected the X handles of each Prolific participant’s followers, resulting in a dataset of a total of n = 1,670,127 users. Finally, the authors collected the latest 500 tweets of each user that contains an URL to external sources, covering a period up to a year before the date of their latest tweet, producing a total of n = 14,688,374 X tweets. From this dataset we obtained all users with available factuality scores, which resulted in a dataset of 19,596 users. Furthermore, we utilise data on the number of followers of the user, number of people followed by the user, total number of tweets produced by the user, registration dates, and other details, all directly extracted from X.
@@
To measure the credibility of news outlets linked in X tweets, Galeazzi et al. (2023) introduce a metric, referred to here, and in their paper, as Factuality. The quantification of this metric relies on data from Media Bias Fact Check (MBFC), an independent fact-checking organisation that classifies news outlets based on both their reliability and political bias [19]. Each news outlet on MBFC is assigned labels indicating its reliability (Very Low, Low, Mixed, Mostly Factual, High, Very High). In the context of their study, news outlets’ factuality corresponds to MBFC’s reliability label. These metrics are applied individually to each X user (denoted as u), generating a vector, $F_{u}\in\mathbb{R}^{6\times 1}$, representing the percentage of links shared by the user for Factuality categories. In this paper we rely on the average of the scores related to the tweets posted by the same user to assign a unique factuality score to each user, which will serve as the basis of our analysis. We segmented users into low, middle, and high factuality groups based on their respective factuality scores. This involved dividing users into the three factuality categories such that the bottom and top 25%, 30%, or 35% would be the low and high factuality users, respectively, and the remaining middle would be the middle factuality group, resulting in three distinct factuality groups. While the main focus of our paper lies on the results derived from the factuality categorisation taking the top and bottom 30% as low and high factuality users, respectively, we also present findings obtained applying the other two thresholds in the Supplementary Information to validate the robustness of our conclusions.
@@
As the purpose of our study is understanding the behaviour of “human users” [20], we initially curated our datasets to focus on users who are more representative of organic human social media users. I.e., we wanted to exclude accounts from the study that appeared to be either bots, business, political or celebrity accounts. To ensure the selection of regular individuals, we implemented specific criteria as outlined below. These criteria were derived from various sources and were applied iteratively to curate our dataset.
@@
Followed account/follower Ratio: A high proportion of non-reciprocal followings may also indicate bot-like behaviour [23]. Therefore, accounts with followed account/follower ratios exceeding certain values were removed. Based on the stringency of the above introduced criteria we created three distinct groups of datasets: relaxed, middle, and strict. By applying these criteria, we aimed to ensure the selection of regular individuals whose behaviour is more representative of typical social media users for our analysis. The specific values for each strictness category are shown in Table 1.
@@
Top/bottom 25% 4899 4612 3172
Top/bottom 30% 5878 5535 3807
Top/bottom 35% 6858 6457 4441
Total 19596 18450 12690
@@ Acknowledgements
Supplementary Information
@@
Figure 1: Low factuality users tend to have higher follower count than high factuality users. Comparing the distributions of follower count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followers compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets.
@@
Figure 3: Low factuality users tend to have higher tweet count than high factuality users. Comparing the distributions of tweet count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of tweets compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets.
@@
Figure 5: Low factuality users tend to have higher followed account count than high factuality users. Comparing the distributions of followed account count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followed accounts compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets. The peak around 5000 is due to an X policy that limits the number of new followed accounts until the user obtains more followers.
@@
Figure 7: Low factuality users tend to have lower number of days since registration than high factuality users. Comparing the distributions of the number of days since registration between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly lower number of days compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is lower than for high factuality users (blue dotted line) in each of the tested datasets.
@@
Through plotting the interactions of follower count with all other metrics, certain differences in the relationships between these metrics and factuality levels across users with varying follower counts have been observed. Specifically, our analysis reveals that in relation to followed account count, users with fewer followers are less likely to exhibit high factuality as they follow more users, whereas those with a larger follower base show a slightly increased likelihood of high factuality with more followed accounts (Figure 10, panel a).Thus, the effect of followed account count on high factuality is strongly negative among those with smaller follower base, while slightly positive among those with larger follower base. For the low factuality group similar trends are observed for both categories of follower count, where users with both few and many followers are less likely to exhibit low factuality as they increase the number of users they follow. Nevertheless, the effect is less strong among those with fewer followers.
@@
In relation to average daily tweet count our analysis (Figure 10, panel b) shows that as users increase their tweet activity, they are less likely to belong to the high factuality category and more likely to belong to the low factuality category, regardless of their follower count. Nevertheless, among users with fewer followers, the effect is more pronounced in the high factuality group compared to the effect observed among users with a larger number of followers. Conversely, in the low factuality group, this relationship is less pronounced among users with fewer followers.
@@
Regarding the relationship between factuality and days since registration, similar patterns emerge across different follower counts (Figure 10, panel c). We show that as users spend more time registered on the platform, they are more likely to belong to the high factuality category and less likely to belong to the low factuality category, and this is consistent in direction and magnitude across users, regardless of their follower count. However, the positive effect of days since registration on high factuality is stronger among those with lower number of followers compared to those with larger number of followers. Nevertheless, we cannot observe such difference in the effects on low factuality.
@@
In examining average daily tweet count across different levels of followed account count and its association with factuality levels (Figure 10, panel d) we find that an increase in tweet activity is associated with a decreased likelihood of belonging to the high factuality group, regardless of followed account count, but the relationship is stronger among those who follow fewer accounts. In contrast, the relationship between tweet count and low factuality varies based on followed account count with a higher degree. Users, who follow a larger number of accounts demonstrate a decreased likelihood of belonging to the low factuality group as tweeting activity increases. Conversely, among users with who follow fewer accounts, higher tweet counts are associated with a somewhat increased likelihood of belonging to the low factuality group.
Our findings regarding the interaction between days since registration and followed account count (Figure 10, panel e) indicate that regardless of followed account count, users who have been registered on the platform for a longer duration are more likely to be in the high factuality category. Conversely, the relationship between days since registration and low factuality varies based on followed account count. The relationship is consistently negative between days since registration and low factuality, indicating that longer registration periods are associated with a decreased likelihood of low factuality, however, this effect is somewhat more pronounced among users who follow less accounts.
@@
Examining the interaction between days since registration and tweet count in relation to factuality levels reveals consistent trends across users with different average daily tweet counts (Figure 10, panel f). Our analysis demonstrates that as the registration period lengthens, users are more likely to exhibit higher factuality levels, and less likely to exhibit lower factuality levels, irrespective of their tweet activity.
@@
Prediction Accuracy
Accuracy scores obtained from shuffled datasets against those derived from the real data to demonstrate that the accuracy obtained from our regression exceeds that expected by random chance.
Figure 11: Accuracy score of predictions using the four social network metrics on the real dataset (red dotted line) compared to the accuracy scores obtained on the 1000 randomised datasets (green bars).

### v4

@@ Discussion
In this study, we aimed to examine whether simple, low-barrier metrics are systematically associated with differences in users’ likelihood of sharing misinformation on the social media X (Twitter). Drawing upon existing literature and leveraging easily accessible data from X, we tested four social network metrics as possible correlates of user factuality. Our analysis focused on follower count, average daily tweet count, number of users followed, and account age. The results offer several insights into the effectiveness of these metrics as group-level indicators of misinformation sharing tendencies.
Our findings indicate that when considered individually, each of the four metrics—follower count, average daily tweet count, number of users followed, and account age—has a significant relationship with user factuality scores. However, when all metrics were analysed together using multinomial regression, the distinctive power of follower count and followed account count diminished for the separation of low and high factuality, while average daily tweet count and number of users followed remained robust correlates of low and high factuality scores.
@@
Finally, the multinomial regression analysis revealed certain significant interaction effects in the case of follower and followed account count. Among users who follow few accounts, tweeting more is less strongly associated with low factuality than among those who follow many accounts, while the effects on high factuality are positive for users who follow many accounts. Among users who follow many accounts, account age is less strongly associated with low factuality than among those who follow few accounts, while among users who follow few accounts, longer time on the platform is linked to higher factuality, whereas among those who follow many accounts, longer platform use is associated with lower factuality. These cross-effects highlight the context-dependent nature of the relationship between these social network metrics and factuality.
@@
These findings have several important implications. For social media platforms, the results suggest that basic metrics like tweet count and account age could inform efforts to understand which types of users are more likely to share misinformation on average, allowing for targeted interventions such as accuracy prompts or digital literacy campaigns. Policymakers could advocate for the incorporation of these metrics into misinformation mitigation strategies, ensuring a low-cost, scalable approach. Future research should continue to explore additional accessible data points and refine existing metrics to better understand the behavioral tendencies associated with misinformation sharing.
While our approach provides valuable insights into the identification of users likely to share misinformation on X, several limitations must be acknowledged. First, it is important to note that the analysis presented here is not aimed at individual-level prediction. Instead, we aim to identify broad trends in user behavior that may be associated with misinformation sharing, rather than develop a robust predictive model. As such, the overlap in distributions among high- and low-factuality users limits the ability to draw definitive conclusions about individual users’ likelihood of spreading misinformation.
Moreover, the dataset we used originates from a specific population—followers of a sample of users recruited via Prolific who self-reported their X handles in a survey on vaccine hesitancy. While this initial sample may introduce some bias, our analyses were not limited to the survey participants themselves but expanded to include their first-degree followers. This significantly larger group—over 11,000 users—likely offers a broader behavioral snapshot of the general Twitter population, though some residual similarities may remain. However, rather than aiming for full population representativity, the purpose of our dataset is to compare the behavior of users who are more prone to share content from unreliable sources to those who are more likely to share reliable ones.
Importantly, our decision to use this dataset was guided by the goal of studying regular users. Many available public datasets that support factuality scoring rely on keyword-based data collection, which tends to over-represent highly active or politically engaged accounts. These are often not reflective of typical user behavior. In contrast, our dataset—filtered to exclude bots, verified users, and extreme activity levels—better represents everyday users and their misinformation-related behaviors. As such, while not perfectly representative, it provides an appropriate test case for our research goals.
Beyond sampling considerations, another limitation of our method is the exclusion of users who share domains without Media Bias/Fact Check (MBFC) ratings. This decision restricts the scope of our analysis, as it may not be fully representative of the broader user base, particularly those sharing content from sources that are not evaluated by MBFC. We collected the latest 500 tweets containing URLs for approximately 1.6 million users. Out of these, 240,915 users had ever shared a URL, and among them, only 59,610 had shared at least one URL that could be classified using MBFC ratings. We report the proportion of such users excluded during our filtering process in the Methods section to provide transparency on the scale of this limitation. Future research could explore the implications of including such users to better understand the full spectrum of misinformation-sharing behavior.
Beyond limitations related to sampling and content coverage, the generalizability of our findings is influenced by the platform-specific nature of user behavior. X, a highly public, text-based platform, is distinct from other social media platforms such as Facebook, which emphasizes private networks and long-form content. The social network metrics we use to identify misinformation spreaders on X may not be as effective on platforms where user engagement patterns differ significantly. Therefore, while our findings offer insights specific to X, they may not fully translate to other platforms with different interaction styles.
Network structure variations between platforms also present challenges. X operates on a directed follower-following model, where information flows asymmetrically, whereas platforms such as Facebook and LinkedIn use mutual connections. This structural difference may result in variations in the usefulness of network-based features like follower count and following count when applied to other platforms with more reciprocal interactions.
The impact of content moderation policies further complicates the application of our approach across platforms. Each platform has unique strategies for content distribution and moderation—such as visibility algorithms, labeling practices, and content removal—that shape the dynamics of misinformation spread. These differences may affect how observable and measurable misinformation sharing is, and thus influence the performance of detection methods like ours.
Despite these limitations, our methodological framework, which leverages social network metrics for early identification of misinformation spreaders, holds promise for adaptation to other platforms. However, careful consideration of platform-specific differences is necessary, and future work should explore these variations to determine whether similar patterns of behavior emerge across different social media environments.
In conclusion, this study demonstrates that basic, easily accessible social network metrics, particularly tweet count and account age, can help characterize differences between low and high factuality users. While the studied metrics may offer a scalable and efficient way to explore behavioral patterns associated with groups of users more likely to share misinformation, they should be interpreted with caution and in context. Their use may be particularly helpful for informing broader strategies aimed at understanding or addressing misinformation dynamics, offering a foundation for cost-effective interventions to curb the spread of misinformation on social media platforms. However, their usefulness may vary across platforms due to differences in user behavior, content moderation policies, and network structures.
@@ Methods
Our study builds upon the data used in the studies by Rathje et al. [22], which was approved by the University of Cambridge Psychology Research Ethics Committee (PRE.2020.144). All participants in the study provided informed consent before answering surveys and providing their Twitter usernames [22]. This dataset was selected in part due to its unique ability to capture regular, human user behavior at scale. Most public misinformation datasets rely on keyword-based collection methods that tend to over-represent highly active or partisan users. Our focus on followers of original survey participants allowed us to build a larger, more behaviorally representative dataset for this purpose.
The study collected 463 survey responses through the online recruitment platform Prolific. Starting from those usernames, the publicly accessible followers of those accounts were collected, resulting in a dataset of n = 1,670,127 users. Such data includes information about the number of followers and following, as well as the number of content posted and the subscription date. For each user, the latest 500 tweets containing a URL to external sources, posted within one year of the most recent published content, were collected. In this study, we utilize users’ data, including the number of followers of the account, the number of people followed by the account, the total number of tweets produced by the account, registration dates, and their tweets containing a URL, all directly extracted using the official Twitter API. All methods in this study were performed in accordance with the relevant guidelines and regulations.
@@
The reliability of users’ diets was measured using a metric called Factuality, based on the domains of the content they published. The quantification of this metric relies on data from Media Bias Fact Check (MBFC), an independent fact-checking organisation that classifies news outlets based on both their reliability and political bias [23]. Each news outlet on MBFC is assigned labels indicating its reliability (Very Low, Low, Mixed, Mostly Factual, High, Very High). News outlets’ factuality corresponds to MBFC’s reliability label. These metrics have been applied individually to each X user (denoted as u), generating a vector, $F_{u}\in\mathbb{R}^{6\times 1}$, representing the percentage of links shared by the user for Factuality categories. We assign each user a unique factuality score, calculated as the average of the factuality scores associated with their posted tweets. To assign factuality scores, we initially collected up to the latest 500 tweets containing URLs, posted within one year of the most recent published content, for the approximately 1.6 million users. Out of these, only 240,915 users had ever shared a URL, and among them, only 59,610 had shared at least one domain we were able to classify using MBFC ratings. We segmented users into low, middle, and high factuality groups based on their respective factuality scores. This involved dividing users into the three factuality categories such that the bottom and top 25%, 30%, or 35% would be the low and high factuality users, respectively, and the remaining middle would be the middle factuality group, resulting in three distinct factuality groups. Our primary goal with this approach is to identify the most prominent misinformation spreaders and characterize their behaviors, rather than provide a continuous score for all users. While continuous scores may offer more granularity, prior work has shown that misinformation sharing on social media follows a highly skewed distribution, with a small fraction of users responsible for the majority of spread [24]. Therefore, segmenting users allows us to better focus on actionable insights relevant to platform moderation and intervention strategies. Although our segmentation approach may appear to contrast with literature highlighting power-law distributions in misinformation sharing, it serves a different purpose: to provide a broad behavioral overview rather than pinpoint only extreme cases. Grouping users in the bottom 30% (or 25%, 35%) allows us to capture general patterns among users with lower factuality. While the main focus of our paper lies on the results derived from the factuality categorisation taking the top and bottom 30% as low and high factuality users, respectively, we also present findings obtained applying the other two thresholds in the Supplementary Information to validate the robustness of our conclusions.
@@
As the purpose of our study is understanding the behaviour of “human users” [25], we initially curated our datasets to focus on users who are more representative of organic human social media users. I.e., we wanted to exclude accounts from the study that appeared to be either business, political or celebrity accounts, or bots. To ensure the selection of regular individuals, we implemented specific criteria as outlined below. These criteria were derived from various sources and were applied iteratively to curate our dataset.
@@
Followed account/follower Ratio: A high proportion of non-reciprocal followings may also indicate bot-like behaviour [28]. Therefore, accounts with followed account/follower ratios exceeding certain values were removed. Finally, to filter for bots, we applied an additional filtering step using Botometer [29], specifically removing users with a Botometer score above 0.4, following the “human threshold” suggested by [30].
Based on the stringency of the above introduced criteria we created three distinct groups of datasets: relaxed, middle, and strict. By applying these criteria, we aimed to ensure the selection of regular individuals whose behaviour is more representative of typical social media users for our analysis. The specific values for each strictness category are shown in Table 1.
To characterize the groups resulting from our filtering criteria, we present the distribution of Factuality scores in Figure 1 in the Supplementary Information, illustrating the differences between the selected groups.
In addition to providing the distributions, we briefly characterize the groups based on their factuality scores. As shown in Supplementary Figure 1, low factuality users are predominantly characterized by scores below 60, middle factuality users span scores between 60 and 70, and high factuality users are characterized by scores above 70, with a strong concentration around 80. We note that the observed peaks in the distributions arise from the underlying labeling procedure: factuality scores are assigned based on categorical source credibility ratings, leading to discrete jumps at 0, 20, 60, 80, and 100. As many users’ sharing histories contain sources with identical credibility ratings, sharp spikes are observed in the plotted distributions at scores 40, 60, and 80.
@@
Top/bottom 25% 4492 4218 2878
Top/bottom 30% 5427 5113 3643
Top/bottom 35% 6166 5806 4072
Total 17616 16586 11503
@@ Acknowledgements
Supplementary Information for “Easy-access online social media metrics are associated with misinformation sharing activity”
Júlia Számely, Alessandro Galeazzi, Júlia Koltai, and Elisa Omodei
@@
Distributions of factuality scores
Figure 1: Distribution of factuality scores across the nine groups, defined by varying levels of regular user and bot filtering criteria (Relaxed, Middle, Strict) and low/high factuality group cutoffs (35/35, 30/30, 25/25).
@@
Figure 2: Low factuality users tend to have higher follower count than high factuality users. Comparing the distributions of follower count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followers compared to high factuality users (p-value << 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets.
@@
Figure 4: Low factuality users tend to have higher tweet count than high factuality users. Comparing the distributions of tweet count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of tweets compared to high factuality users (p-value << 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets.
@@
Figure 6: Low factuality users tend to have higher followed account count than high factuality users. Comparing the distributions of followed account count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followed accounts compared to high factuality users (p-value << 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets. The peak around 5000 is due to an X policy that limits the number of new followed accounts until the user obtains more followers.
@@
Figure 8: Low factuality users tend to have lower number of days since registration than high factuality users. Comparing the distributions of the number of days since registration between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly lower number of days compared to high factuality users (p-value << 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is lower than for high factuality users (blue dotted line) in each of the tested datasets.
@@
Through plotting the interactions of follower count with all other metrics, certain differences in the relationships between these metrics and factuality levels across users with varying follower counts have been observed. Specifically, our analysis reveals that in case of users with fewer followers the effect of the followed account count diminishes on high factuality, whereas those with a larger follower base show an increased likelihood of high factuality with more followed accounts (Figure 11, panel a). Thus, the effect of followed account count on high factuality is nonexistent among those with smaller follower base, while positive among those with larger follower base. For users in the low factuality group, we observe a similar trend across both follower count categories: as users follow more accounts, they become less likely to belong to the low factuality group. This effect is slightly stronger for users with a higher follower count. Among those with fewer followed accounts, popular users are more likely to have low factuality, while in case of those who follow many accounts, there is no significant difference in their likelihood of low factuality by popularity.
@@
In relation to average daily tweet count our analysis (Figure 11, panel b) shows that as users increase their tweet activity, they are less likely to belong to the high factuality category and more likely to belong to the low factuality category, regardless of their follower count. Nevertheless, in the low factuality group, this relationship is less pronounced among users with fewer followers.
@@
Regarding the relationship between factuality and days since registration (Figure 11, panel c) we show that as users spend more time registered on the platform, they are more likely to belong to the high factuality category and less likely to belong to the low factuality category, and this is consistent in direction and magnitude across users, regardless of their follower count. However, the positive effect on high factuality is slightly stronger among users with higher follower counts. For the low factuality group, we observe a negative relationship with days since registration, with a marginally stronger effect among those with more followers. It shows that among younger users popular ones are more likely to have lower factuality compared to less popular ones. However, among older users there is no difference in the likelihood of factuality between more and less popular users.
@@
In examining average daily tweet count across different levels of followed account count and its association with factuality levels (Figure 11, panel d) we find that among those with few accounts followed the average tweet count does not affect the likelihood of belonging to the high factuality group. However, an increase in tweet activity is associated with an increased likelihood of belonging to the high factuality group among those who follow more accounts. For the low factuality group, tweet count is negatively associated with low factuality and this effect is stronger among those who follow more accounts. There is no significant difference in the likelihood of belonging to the low factuality group among those who tweet infrequently, however, among those who tweet frequently, users following few accounts are way more likely to have low factuality than users who follow more accounts. Figure 11, panel d corresponds to Figure 6, panel a in the main text.
@@
Examining the interaction between days since registration and tweet count in relation to factuality levels reveals consistent trends across users with different average daily tweet counts (Figure 11, panel e). Our analysis demonstrates that as the registration period lengthens, users are more likely to exhibit higher factuality levels, and less likely to exhibit lower factuality levels, irrespective of their tweet activity.
Our findings regarding the interaction between days since registration and followed account count (Figure 11, panel f) indicate that among users who follow fewer accounts, those who have been registered on the platform for a longer duration are more likely to be in the high factuality category, while among those who follow more accounts, the trend is the opposite: the longer they are on the platform, the less likely that they belong to the high factuality group. The relationship between days since registration and low factuality does not vary based on followed account count. The relationship is consistently negative between days since registration and low factuality, indicating that longer registration periods are associated with a decreased likelihood of low factuality, however, this effect is somewhat more pronounced among users who follow fewer accounts. Figure 11, panel f corresponds to Figure 6, panel b in the main text.
@@
Pseudo-R (McFaddens R)
Pseudo-R (McFaddens R) obtained from shuffled datasets against those derived from the real data to demonstrate that the Pseudo-R obtained from our regression exceeds that expected by random chance.
Figure 12: Pseudo-R (McFaddens R) using the four social network metrics on the real dataset (red dotted line) compared to the accuracy scores obtained on the 1000 randomised datasets (blue bars).